Two draws are made at random without replacement from the digits {1,2,3,4}. Let X1 be the first digit drawn and X2 the second.

Let M=max(X1,X2) and S=X1+X2.

a) Find ?(?).

b) Make a joint distribution table for M and S.

c) Use the table in Part b to find the distribution of M.

d) Find ?(?).
Answer #1

Two draws are made at random without replacement from the digits {1,2,3,4}.

Let X1 be the first digit drawn and X2 the second.

So, we make the combination as (X1,X2)

{(1, 2), (1,3), (1,4), (2,1), (2,3), (2,4), (3,1), (3,2), (3,4), (4,1), (4,2), (4,3)}

I had a 1/4 probability of choosing that one and then 1/3 probability of choosing that second.

For each combination, the probability is 1/12

M=max(X1,X2) = {2,3,4}

M Probability
2 1/6
3 1/3
4 1/2

S=X1+X2 = {3,4,5,6,7}

S Probability
3 1/6
4 1/6
5 1/3
6 1/6
7 1/6

a)

b)

M
S 2 3 4 Total
3 (1/6) 0 0 1/6
4 0 (1/6) 0 1/6
5 0 (1/6) (1/6) 1/3
6 0 0 (1/6) 1/6
7 0 0 (1/6) 1/6
Total 1/6 1/3 1/2 1

c)

M Probability
2 1/6
3 1/3
4 1/2

d)